# Swivuriso: The South African Next Voices Multilingual Speech Dataset

### v1

@@ 1 Introduction
Automatic speech recognition (ASR) has advanced significantly, mainly due to the emergence of deep learning methods and the increasing availability of extensive multilingual datasets [Dimah2024, KHEDDAR2024102422, Tjandra2023, Zeng2022, pmlr-v202-radford23a]. However, many African languages continue to be underrepresented in these advancements [Nakatumba-Nabende2024, elamin2023multilingual]. Limited availability of quality speech datasets for these languages restricts the creation and deployment of accurate and inclusive ASR systems [Etienne2014, BADENHORST2022107860].
Current ASR data sets often do not reflect the rich linguistic and cultural diversity of the African continent [dossou2023advancing, chang2024self]. They are generally limited in scope [mak2024], focusing predominantly on general-purpose speech. Moreover, these datasets are typically composed of scripted recordings, which limits the ability of ASR systems to handle spontaneous speech and multilingual interactions that are common in many African communities [BiswasYWWN20].
@@ 2.1 Coverage and Diversity
More recent datasets include Mozilla’s Common Voice, which provides crowd-sourced, scripted speech in six South African languages, distributed under a permissive CC-0 license to encourage broad reuse. FLEURS includes approximately 12 hours of data per language in four South African languages, designed primarily for benchmarking and evaluation. Vuk’uzenzele(ViXSD)[rajab2025esethu] is a community-led ASR dataset, released under the Esethu licensing framework, emphasizing community ownership and ethical commercialization. While smaller in scale, it highlights local participation in resource creation. AfriSpeech‑200 [Afrispeech-200] includes English speech with South African accents and healthcare domain coverage, useful for accent-aware ASR modeling.
@@ 2.2 Ethical and Participatory Data Collection Practices
Collecting speech data in underrepresented language communities demands sensitivity to consent, privacy, cultural norms, and long-term community impact [holton2021people, armentano-oller-etal-2024-becoming]. Prior work in projects such as Common Voice and ALFFA [besacier2015speech] has demonstrated the importance of embedding ethical and participatory practices throughout the data lifecycle. A recent case study on the Catalan edition of Common Voice highlights the need for clear consent procedures, privacy safeguards, and inclusive participation [armentano-oller-etal-2024-becoming]. Similarly, the ALFFA project emphasised community involvement by working with local experts and field linguists during data collection [besacier2015speech].
@@ 2.3 Licensing and Accessibility
Licensing and accessibility are essential for ensuring the broad usability of datasets, particularly for low-resource languages. Open licenses–such as Creative Commons–have been widely recognized for enabling broad reuse, especially where data collection is limited or financially inaccessible [Pratap2020MLSAL]. Projects like Common Voice illustrate how permisssive licensing can support equitable access and accelerate language technology development for underrepresented languages [ardila-etal-2020-common].
In contrast, access-restricted datasets, while valuable, remain difficult to obtain for many researchers–particularly those based in low- and middle-income countries (LMICs). Yet, concerns around dataset accessibility are only part of a broader conversation. Increasingly, the focus is shifting toward ethical and equitable data governance, especially in contexts where language data is sourced from communities that have historically had limited control over how it is used. This has prompted interest in alternative licensing models that prioritize responsible stewardship, local agency, and benefit sharing [OkorieOminoLicensing, rajab2025esethu].
@@ 3.2 Dataset Creation Process
Remuneration was stratified by language to reflect engagement needs [de2017speech]. This workflow helped ensure that the final dataset reflects high technical quality, diverse speaker representation, and strong ethical oversight.
@@ 3.3.2 Scripted Prompts
  • Pula/Imvula Magazine [grainsa_pula_imvula], a monthly publication of Grain SA [^2] https://www.grainsa.co.za providing agricultural information in five South African languages (English, Sesotho, Setswana, isiZulu and isiXhosa). This data was made available through South African Centre for Digital Language Resources (SADiLaR) [^3] https://sadilar.org/.
  • Vukuzenzele [lastrucci2023preparingvukuzenzelezagovmultilingualsouth], a government publication that offers information on public services.
@@
  • African Wordnet [sonja_marissa_wordnet], a vital lexical database to ensure vocabulary accuracy and support the development of various language technologies in South African languages.
  • Translated goatkeepers manuals (ARC animal health) [vatta2007Goatkeepers], specialized content on animal health and agriculture.
@@
To standardize the content and facilitate thematic analysis, two steps were performed: First, the entire text was translated into English using NLLB-200 [nllbteam2022languageleftbehindscaling]. Second, translated text files were categorized into themes (e.g., agriculture, healthcare, general) using OpenAI’s GPT-4 [openai2023gpt4].
@@
The text was segmented into meaningful sentences (prompts) using BlingFire [microsoft2021blingfire]. This ensured accurate segmentation by extracting discrete sentences, enforcing proper punctuation at each sentence end, and merging very short sentences to avoid fragmentation.
Finally, Named Entity Recognition (NER) and Personal Information Anonymization were applied. Specifically, Personally Identifiable Information (PII) was flagged or removed using Microsoft Presidio [microsoftPresidio], which identifies various PII types.
@@ 3.4 Dataset Statistics
The above deterministic notation captures the key statistics of the Swivuriso dataset and is aligned with standard practice evident in many contemporary state-of-the art multilingual datasets  [Pratap2020MLSAL, conneau2023fleurs, galvez2021people]. It accounts for the number of clips, clip durations, speaker contributions, domain coverage, and word counts for both scripted and unscripted recordings.
@@ 4.2 Whisper-large-v3-Turbo Monolingual Baselines
Training Configuration: All training was conducted for 20002000 steps with an effective batch size of 3232 and a learning rate of $1\times 10^{-4}$.
@@ 4.4 Effect of Speech hours in Training data
Training Configuration: All training was 1 epoch with an effective batch size of 6464 and a learning rate of $1\times 10^{-4}$.
@@ 5 Related Work
There has been substantial work done previously by many scholars in the art of collecting and designing world class datasets especially in the field of Automatic Speech Recognition where the requirements for precision and elimination of anomalies are very important [junczyk2024survey, Afrispeech-200, kahn2020libri, yeroyan2024enabling, virkkunen2023finnish, kolobov2021mediaspeech]. In the context of low resource languages globally the development has been slow particularly for African languages. We take note of the ones that has helped enlighten us in this journey Abdulmumin et. el proposed a study [abdulmumin2024correcting] on the FLORES dataset which helped correct FLORES to better evaluate four African low resource languages. This work involved a meticulous review and correction of the dataset’s dev and devtest splits for Hausa, Northern Sotho (Sepedi), Xitsonga, and isiZulu. In this study, native speakers carefully identified and fixed inconsistencies and inaccuracies that could impact the reliability of downstream NLP tasks, especially machine translation. The work further demonstrated that their corrections improved the linguistic accuracy and overall quality of the dataset, supported by both qualitative assessment and statistical analysis of the changes. This correction effort highlights the importance of native speaker involvement to ensure linguistic and cultural relevance in low-resource language datasets. We have heavily involved native speakers in our dataset to strengthen its character and overall improve the linguistic accuracy and the overall quality of the dataset as recommended by [abdulmumin2024correcting]. Another interesting study which helped guide our process in terms of sentiment analysis across a range of domains is the study proposed by authors in [ekbal2022hindimd]. The study identified that many sentiment analysis benchmark datasets for high resource languages have been made available however, for low-resource languages this is not the case especially for Indian languages like Hindi. The study addressed the challenges of sentiment analysis in Hindi particularly addressing socially relevant domains by introducing a balanced corpus annotated with the sentiment classes, viz. positive, negative and neutral.
@@ 6 Licensing and Accessibility
Dataset licensing is fundamental for ensuring both broad usability and ethical access, especially for low-resource languages. Standard open licenses, such as Creative Commons (e.g., $C$ BY 4.0), have made large-scale reuse possible, with initiatives like Mozilla Common Voice demonstrating the value of permissive licensing for language technology development. However, access restrictions to resources for AI development continue to impede many researchers in low- and middle-income countries, pointing to broader concerns about equitable data governance and local community agency [OkorieOminoLicensing].
For the Swivuriso project, we adopted the Creative Commons Attribution 4.0 International ($C$ BY 4.0) license to promote open access while embedding ethical accountability. Swivuriso is released on Zenodo and Hugging Face, with persistent DOIs and comprehensive documentation. Each release includes audio, aligned transcripts, and detailed metadata on demographic, linguistic, and accessibility factors. We also provide an ethical use statement, restricting uses such as voice cloning to protect speaker privacy.
The $C$ BY 4.0 license is highly permissive: users can share, adapt, and build upon the dataset for any purpose, including commercially, provided they attribute the creators and indicate any modifications. The license prohibits additional legal or technical restrictions, covers database rights, and includes a “cure period” for addressing license violations, all designed to encourage legal certainty and global scientific collaboration.
However, while open licenses are essential for maximising data utility, they often fail to address issues like local agency or protection of Indigenous knowledge. The uniform application of standard licenses can unintentionally undermine community ownership, especially for cultural materials or community-sourced data. This limitation is driving a shift toward more context-sensitive frameworks that prioritize stewardship, benefit-sharing, and community control. This an approach that is gaining ground in African data governance [rajab2025esethu, OkorieOminoLicensing].
Recent licensing models such as the Nwulite Obodo Open Data License (NOODL) [okorie2025s, okorie2025addressing] and the Esethu License [rajab2025esethu] illustrate this paradigm shift. NOODL, for example, differentiates between users in Africa or developing countries and those outside these regions, enabling tailored benefit-sharing and requiring adaptations to be shared within the region. Esethu reinvests licensing revenue directly into expanding resources and supporting local employment, maintaining both community agency and resource sustainability.
@@
Our choice to use $C$ BY 4.0 for Swivuriso was rooted in a desire for openness, but it also underscored the shortcomings of standard licenses for datasets from African and low-resource languages. Emerging frameworks like NOODL and Esethu highlight the field’s move towards more ethical, equitable governance. For dataset publications, it is no longer sufficient to declare a standard open license; authors should also explain their approach to ethical, just, and community-empowering data stewardship.

### v2

@@ 1 Introduction
Automatic speech recognition (ASR) has advanced significantly, mainly due to the emergence of deep learning methods and the increasing availability of extensive multilingual datasets [2, 21, 38, 42, 34]. However, many African languages continue to be underrepresented in these advancements [27, 15]. Limited availability of quality speech datasets for these languages restricts the creation and deployment of accurate and inclusive ASR systems [6, 5].
Current ASR data sets often do not reflect the rich linguistic and cultural diversity of the African continent [13, 10]. They are generally limited in scope [24], focusing predominantly on general-purpose speech. Moreover, these datasets are typically composed of scripted recordings, which limits the ability of ASR systems to handle spontaneous speech and multilingual interactions that are common in many African communities [8].
@@ 2.1 Coverage and Diversity
More recent datasets include Mozilla’s Common Voice, which provides crowd-sourced, scripted speech in six South African languages, distributed under a permissive CC-0 license to encourage broad reuse. FLEURS includes approximately 12 hours of data per language in four South African languages, designed primarily for benchmarking and evaluation. Vuk’uzenzele(ViXSD)[35] is a community-led ASR dataset, released under the Esethu licensing framework, emphasizing community ownership and ethical commercialization. While smaller in scale, it highlights local participation in resource creation. AfriSpeech‑200 [31] includes English speech with South African accents and healthcare domain coverage, useful for accent-aware ASR modeling.
@@ 2.2 Ethical and Participatory Data Collection Practices
Collecting speech data in underrepresented language communities demands sensitivity to consent, privacy, cultural norms, and long-term community impact [18, 4]. Prior work in projects such as Common Voice and ALFFA [7] has demonstrated the importance of embedding ethical and participatory practices throughout the data lifecycle. A recent case study on the Catalan edition of Common Voice highlights the need for clear consent procedures, privacy safeguards, and inclusive participation [4]. Similarly, the ALFFA project emphasised community involvement by working with local experts and field linguists during data collection [7].
@@ 2.3 Licensing and Accessibility
Licensing and accessibility are essential for ensuring the broad usability of datasets, particularly for low-resource languages. Open licenses–such as Creative Commons–have been widely recognized for enabling broad reuse, especially where data collection is limited or financially inaccessible [33]. Projects like Common Voice illustrate how permisssive licensing can support equitable access and accelerate language technology development for underrepresented languages [3].
In contrast, access-restricted datasets, while valuable, remain difficult to obtain for many researchers–particularly those based in low- and middle-income countries (LMICs). Yet, concerns around dataset accessibility are only part of a broader conversation. Increasingly, the focus is shifting toward ethical and equitable data governance, especially in contexts where language data is sourced from communities that have historically had limited control over how it is used. This has prompted interest in alternative licensing models that prioritize responsible stewardship, local agency, and benefit sharing [28, 35].
@@ 3.2 Dataset Creation Process
Remuneration was stratified by language to reflect engagement needs [12]. This workflow helped ensure that the final dataset reflects high technical quality, diverse speaker representation, and strong ethical oversight.
@@ 3.3.2 Scripted Prompts
  • Pula/Imvula Magazine [36], a monthly publication of Grain SA [^2] https://www.grainsa.co.za providing agricultural information in five South African languages (English, Sesotho, Setswana, isiZulu and isiXhosa). This data was made available through South African Centre for Digital Language Resources (SADiLaR) [^3] https://sadilar.org/.
  • Vukuzenzele [23], a government publication that offers information on public services.
@@
  • African Wordnet [9], a vital lexical database to ensure vocabulary accuracy and support the development of various language technologies in South African languages.
  • Translated goatkeepers manuals (ARC animal health) [39], specialized content on animal health and agriculture.
@@
To standardize the content and facilitate thematic analysis, two steps were performed: First, the entire text was translated into English using NLLB-200 [37]. Second, translated text files were categorized into themes (e.g., agriculture, healthcare, general) using OpenAI’s GPT-4 [32].
@@
The text was segmented into meaningful sentences (prompts) using BlingFire [25]. This ensured accurate segmentation by extracting discrete sentences, enforcing proper punctuation at each sentence end, and merging very short sentences to avoid fragmentation.
Finally, Named Entity Recognition (NER) and Personal Information Anonymization were applied. Specifically, Personally Identifiable Information (PII) was flagged or removed using Microsoft Presidio [26], which identifies various PII types.
@@ 3.4 Dataset Statistics
The above deterministic notation captures the key statistics of the Swivuriso dataset and is aligned with standard practice evident in many contemporary state-of-the art multilingual datasets  [33, 11, 16]. It accounts for the number of clips, clip durations, speaker contributions, domain coverage, and word counts for both scripted and unscripted recordings.
@@ 4.2 Whisper-large-v3-Turbo Monolingual Baselines
Training Configuration: All training was conducted for $2000$ steps with an effective batch size of $32$ and a learning rate of $1\times 10^{-4}$.
@@ 4.4 Effect of Speech hours in Training data
Training Configuration: All training was 1 epoch with an effective batch size of $64$ and a learning rate of $1\times 10^{-4}$.
@@ 5 Related Work
There has been substantial work done previously by many scholars in the art of collecting and designing world class datasets especially in the field of Automatic Speech Recognition where the requirements for precision and elimination of anomalies are very important [19, 31, 20, 41, 40, 22]. In the context of low resource languages globally the development has been slow particularly for African languages. We take note of the ones that has helped enlighten us in this journey Abdulmumin et. el proposed a study [1] on the FLORES dataset which helped correct FLORES to better evaluate four African low resource languages. This work involved a meticulous review and correction of the dataset’s dev and devtest splits for Hausa, Northern Sotho (Sepedi), Xitsonga, and isiZulu. In this study, native speakers carefully identified and fixed inconsistencies and inaccuracies that could impact the reliability of downstream NLP tasks, especially machine translation. The work further demonstrated that their corrections improved the linguistic accuracy and overall quality of the dataset, supported by both qualitative assessment and statistical analysis of the changes. This correction effort highlights the importance of native speaker involvement to ensure linguistic and cultural relevance in low-resource language datasets. We have heavily involved native speakers in our dataset to strengthen its character and overall improve the linguistic accuracy and the overall quality of the dataset as recommended by [1]. Another interesting study which helped guide our process in terms of sentiment analysis across a range of domains is the study proposed by authors in [14]. The study identified that many sentiment analysis benchmark datasets for high resource languages have been made available however, for low-resource languages this is not the case especially for Indian languages like Hindi. The study addressed the challenges of sentiment analysis in Hindi particularly addressing socially relevant domains by introducing a balanced corpus annotated with the sentiment classes, viz. positive, negative and neutral.
@@ 6 Licensing and Accessibility
Dataset licensing is fundamental for ensuring both broad usability and ethical access, especially for low-resource languages. Standard open licenses, such as Creative Commons (e.g., CC BY 4.0), have made large-scale reuse possible, with initiatives like Mozilla Common Voice demonstrating the value of permissive licensing for language technology development. However, access restrictions to resources for AI development continue to impede many researchers in low- and middle-income countries, pointing to broader concerns about equitable data governance and local community agency [28].
For the Swivuriso project, we adopted the Creative Commons Attribution 4.0 International (CC BY 4.0) license to promote open access while embedding ethical accountability. Swivuriso is released on Zenodo and Hugging Face, with persistent DOIs and comprehensive documentation. Each release includes audio, aligned transcripts, and detailed metadata on demographic, linguistic, and accessibility factors. We also provide an ethical use statement, restricting uses such as voice cloning to protect speaker privacy.
The CC BY 4.0 license is highly permissive: users can share, adapt, and build upon the dataset for any purpose, including commercially, provided they attribute the creators and indicate any modifications. The license prohibits additional legal or technical restrictions, covers database rights, and includes a “cure period” for addressing license violations, all designed to encourage legal certainty and global scientific collaboration.
However, while open licenses are essential for maximising data utility, they often fail to address issues like local agency or protection of Indigenous knowledge. The uniform application of standard licenses can unintentionally undermine community ownership, especially for cultural materials or community-sourced data. This limitation is driving a shift toward more context-sensitive frameworks that prioritize stewardship, benefit-sharing, and community control. This an approach that is gaining ground in African data governance [35, 28].
Recent licensing models such as the Nwulite Obodo Open Data License (NOODL) [30, 29] and the Esethu License [35] illustrate this paradigm shift. NOODL, for example, differentiates between users in Africa or developing countries and those outside these regions, enabling tailored benefit-sharing and requiring adaptations to be shared within the region. Esethu reinvests licensing revenue directly into expanding resources and supporting local employment, maintaining both community agency and resource sustainability.
@@
Our choice to use CC BY 4.0 for Swivuriso was rooted in a desire for openness, but it also underscored the shortcomings of standard licenses for datasets from African and low-resource languages. Emerging frameworks like NOODL and Esethu highlight the field’s move towards more ethical, equitable governance. For dataset publications, it is no longer sufficient to declare a standard open license; authors should also explain their approach to ethical, just, and community-empowering data stewardship.